OpenStudy (anonymous):

Limit/Taylor series question:

OpenStudy (anonymous):

\[\lim_{x \rightarrow 0}\large \frac{\cos(x)-1+\frac{x^{2}}{2}}{9x^{4}}\]

OpenStudy (anonymous):

the first three terms of the taylor series of cos(x) are\[1-\frac{x^{2}}{2}+\frac{x^{4}}{24}\] so if you sub those in for cos(x) then everything cancels and you're left with \[\frac{1}{216}\]

OpenStudy (anonymous):

looks good to me

OpenStudy (anonymous):

my questions is why would i arbitrarily stop after the third term....if I continue with the taylor series then when i take the limit to zero i will be dividing by zero

OpenStudy (anonymous):

the rest is automatically zero

OpenStudy (anonymous):

for example next term is taylor series is has a \(x^6\) when you divide by \(x^4\) you get \(x^2\) and that clearly goes to zero

OpenStudy (anonymous):

so all the rest of the terms have positive exponents

OpenStudy (anonymous):

\[\Large \frac{-\frac{x^{6}}{24}}{9x^{4}}\] will be undefined at x=0

OpenStudy (anonymous):

no, it will be \(-x^2\)

OpenStudy (anonymous):

oooooooooooooooooooooooh.....i see....rookie mistake....cancel the x's and the limit wont be a problem

OpenStudy (anonymous):

laws of exponents still work in this situation right?

OpenStudy (anonymous):

right!!

OpenStudy (anonymous):

thanks :)